OpenStudy (study312):

How do you find the square root of a number without a calculator? Whoever can show me how gets a medal

OpenStudy (anonymous):

yes there is a way!

OpenStudy (addicted2anime):

yes, Babylonian method, dont ask me how, lol

OpenStudy (study312):

can one of you tell me please?

OpenStudy (addicted2anime):

Im not sure how, sorry.

OpenStudy (addicted2anime):

you could look it up thou.

OpenStudy (study312):

whoever can show me how gets a medal

OpenStudy (kayders1997):

Well if you have a perfect square than you can 1 4 9 16 25 etc... or if your want to simplify square root of 20 you could

OpenStudy (kayders1997):

What number are you looking at?

OpenStudy (study312):

25

OpenStudy (kayders1997):

So than if it this way the opposite is squaring so x^2=25 what would x be?

jhonyy9 (jhonyy9):

for example sqrt(3617) = ?

OpenStudy (study312):

well 25 ÷ 2 = 12.5, x = 12.5. idk how to do x to the of 2 exactly

OpenStudy (kayders1997):

X=sqrt25 square both sides x=(sqrt25)2 so x^2=25 and because your squaring with a square root the square root does away this is saying x times x=25 so what two numbers that are the same is 25?

OpenStudy (study312):

5 • 5 = 25

OpenStudy (kayders1997):

Yes so your answer is 5

OpenStudy (study312):

oh ok, thx.

OpenStudy (kayders1997):

Does this make sense?

OpenStudy (study312):

yes thank you

OpenStudy (kayders1997):

Good and you're welcome :)

OpenStudy (study312):

now I don't have to keep going to google everytime I get square root questions.

OpenStudy (kayders1997):

Yay :)